OpenStudy (anonymous):

What is the multiplicative inverse of -1/9? (help me understand this question)

OpenStudy (photon336):

\[n * \frac{ 1 }{ n } = \frac{ n }{ n } = 1\]

OpenStudy (photon336):

this is generally what the multiplicative inverse is

OpenStudy (photon336):

\[\frac{ -1 }{ 9 }*(-9/1) = (\frac{ -9 }{ -9 }) = 1\]

OpenStudy (photon336):

@MissLearning please tell me if this makes sense or not

OpenStudy (anonymous):

It does :) ! im sorry for the late reply Im watching a video about to !! Thank you !!

OpenStudy (photon336):

@MissLearning Can you tell me what the multiplicative inverse is for say (1/3)?